OpenStudy (anonymous):

A period of a simple pendulum is 2.4s. The length of the pendulum is doubled. What is the period T of the longer pendulum

OpenStudy (irishboy123):

\[T = 2\pi \sqrt{\frac{L}{g}}\] nuff said?

OpenStudy (shamim):

For first case T1=2*pi*root(L1/g) 2.4=2*pi*root(L/g).............(1)

OpenStudy (shamim):

For second case T2=2*pi*root(L2/g) T2=2*pi*root(2L/g)..........(2)

OpenStudy (shamim):

Now u hv to divide eqution 2 by equation 1

OpenStudy (shamim):

U will get T2/2.4=root(2) T2=?

OpenStudy (shamim):

Response plz!!!!!!!!